# Dictionary-Learning-Based Data Pruning for System Identification

### v1

@@ Abstract
System identification is normally involved in augmenting time series data by time shifting and nonlinearisation (via polynomial basis), which introduce redundancy both feature-wise and sample-wise. Many research works focus on reducing redundancy feature-wise, while less attention is paid to sample-wise redundancy. This paper proposes a novel data pruning method, called (mini-batch) FastCan, to reduce sample-wise redundancy based on dictionary learning. Time series data is represented by some representative samples, called atoms, via dictionary learning. The useful samples are selected based on their correlation with the atoms. The method is tested on one simulated dataset and two benchmark datasets. The R-squared between the coefficients of models trained on the full and the coefficients of models trained on pruned datasets is adopted to evaluate the performance of data pruning methods. It is found that the proposed method significantly outperforms the random pruning method.
@@ 1 Introduction
To avoid model overfitting and reduce computational complexity, feature selection methods based on orthogonalisation techniques and greedy search are widely applied in constructing NARX models [10, 11, 12]. The orthogonalisation-based method was firstly derived in [10] to efficiently decide which terms should be included in the Nonlinear Autoregressive Moving Average with eXogenous input (NARMAX) models. Then this method was further developed in [11] as orthogonal forward-regression estimators to identify parsimonious models of structure-unknown systems by modifying and augmenting some orthogonal least squares methods. A more comprehensive review of this feature selection idea and its development for system identification can be found in [12].
@@
Similar to feature selection, input selection is also a crucial step in system identification, which can help to improve the identification performance and reduce storage and training costs [13, 14, 12]. While optimal input selection (often referred to as data pruning) is currently an active research topic for computer vision [15, 16, 17] and natural language processing [18, 19], its application to system identification has received less attention, although it remains an intractable task.
@@
This paper proposes a data-pruning method based on dictionary learning to select useful time series samples for constructing the simplest mathematical model to represent a system. Specifically, the reduced polynomial NARX method is applied to the identification of nonlinear dynamic systems as an example. The canonical-correlation-based fast feature selection method introduced in [20] is adopted to find the most important terms that should be included in the NAXR model to achieve the required accuracy. Dictionary learning is then creatively combined with the fast feature selection method based on canonical correlation to find the most useful time series samples to train these terms.
@@
Dictionary learning, initially popular in the signal processing area, is used to learn an overcomplete dictionary composed of signal atoms, where a sparse linear combination of these atoms can generate the original signal [21]. It is currently widely studied for image processing since images usually admit sparse representation [22, 23]. For the proposed method, k-means-based dictionary learning provides an overcomplete dictionary of time terms, which are used as pseudo-labels in the data pruning process.
@@
In this paper, the proposed method, i.e., (mini-batch) FastCan, is compared with the random pruning method. The reduced NARX trained with the full dataset is used as the baseline model. The coefficients learned with the pruned dataset is compared with the baseline model. The more the coefficients are close to the baseline model, the better the pruned dataset is. The procedural of the research in this paper is illustrated in Fig. 1.
@@
The next section introduces the details of the proposed data-pruning method for system identification, followed by numerical case studies to demonstrate the necessity and advantages of the method. Subsequently, case studies on two benchmark datasets for nonlinear system identification are presented, with a discussion on the effect of hyperparamteres in the proposed method. It is worth noting that the proposed method can also be used to select informative samples for general machine-learning tasks.
@@
Figure 1: The schematic diagram of this research. The time series $y(t)$ is generated by the target equation, then time-shifted and nonlinearised to construct a redundant term library. Feature-wise redundancy is represented by a correlation heatmap, while sample-wise redundancy is visualised by a principle component analysis (PCA) plot. Feature selection is applied to the term library to reduce feature-wise redundancy. Dataset A [^1] The colour style used this figure is from bg-mpl-stylesheets [24]., which contains the full set of samples, is used to train the Baseline model. Dataset B, a random subset of Dataset A, is used to train the Random pruned model. Dataset C, consisting of samples selected based on the atoms learned from Dataset A, is used to train the FastCan pruned model. The FastCan pruned dataset generally exhibits less sample-wise redundancy compared to the randomly pruned dataset, as illustrated in the PCA plot, where the data selected by FastCan are more widely spread, while the randomly pruned data tend to be close to each other. To compare the data pruning methods, Dataset A is pruned repeatedly by random selection and FastCan, and the models are then trained on the pruned datasets, respectively. The results show that the coefficients learned from the FastCan pruned data are closer to those of the baseline model, with less variance across repeated tests.
@@ 2 Methodology
This section introduces the dictionary learning-based data pruning method for a discrete system with finite orders.
@@ 2.1 Reduced polynomial NARX
A NARX model is used to simulate a system, which is formulated as
$\begin{split}y(k)=&F[y(k-1),y(k-2),...,y(k-n_{y}),\\ &u(k-d),y(k-d-1),...,y(k-d-n_{u})]+e(k)\end{split}$ (1)
where $y(k)$, $u(k)$ and $e(k)$ are the system output, input and noise sequences, respectively. $n_{y}$ and $n_{u}$ are the maximum terms for the system output and input. $d$ is the time delay, typically set to $d=1$. $F[\cdot]$ is a nonlinear function.
@@ A NARX model is used to simulate a system, which is formulated as
$\begin{split}y(k)=\theta_{0}&+\sum_{i_{1}=1}^{n}f_{i_{1}}\big{(}x_{i_{1}}(k)\big{)}\\ &+\sum_{i_{1}=1}^{n}\sum_{i_{2}=i_{1}}^{n}f_{i_{1}i_{2}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k)\big{)}+\dots+\sum_{i_{1}=1}^{n}\cdots\\ &\sum_{i_{\ell}=i_{\ell-1}}^{n}f_{i_{1}i_{2}\cdots i_{\ell}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k),\dots,x_{i_{\ell}}(k)\big{)}+e(k)\end{split}$ (2)
@@
$f_{i_{1}i_{2}\cdots i_{m}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k),\dots,x_{i_{m}}(k)\big{)}=\theta_{i_{1}i_{2}\cdots i_{m}}\prod_{k=1}^{m}x_{i_{k}}(k),$
$x_{m}(k)=\begin{cases}y(k-m),&1\leq m\leq n_{y},\\ u(k-m+n_{y}),&n_{y}+1\leq m\leq n=n_{y}+n_{u},\end{cases}$
where $1\leq m\leq\ell$, $\theta_{i_{1}i_{2}\cdots i_{m}}$ are model parameters, and $\prod_{k=1}^{m}x_{i_{k}}(k)$ are model terms whose order is not higher than $\ell$.
The total number of model terms in the polynomial NARX model given in equation (2) is $M=(n+\ell)!/[n!\,\ell!]$. It can be seen that the full NARX model can include a large number of terms, increasing the risk of overfitting. In addition, only a subset of these terms is typically important to capture the underlying dynamic relationship [2]. Therefore, the canonical-correlation-based fast feature selection is carried out to find out the significant model terms from $\prod_{k=1}^{m}x_{i_{k}}(k)$ with $y(k)$ as the target. Refer to [20] for details on the specific feature selection steps.
2.2 Mini-batch data pruning
@@
$\mathbf{X}=\begin{pmatrix}x_{1,1}&\dots&x_{1,N}\\ \vdots&\ddots&\vdots\\ x_{p,1}&\dots&x_{p,N}\end{pmatrix}$
where $p$ is the number of the features, that is the selected model terms, and $N$ is the number of time series samples given the selected features.
A dictionary $\mathbf{D}\in\mathbb{R}^{p\times K}$ for the matrix $\mathbf{X}\in\mathbb{R}^{p\times N}$ is obtained by mini-batch k-means clustering [25] and used as the target $\mathbf{Y}$ for the subsequent sample selection step, where $K$ is the number of atoms in the dictionary, and typically $K>p$ for an overcomplete dictionary. Each data sample $\mathbf{x}_{j}\in\mathbb{R}^{p}$ (column of $\mathbf{X}$) can be approximated as a sparse linear combination of the dictionary atoms, given by,
@@
where $\mathbf{a}_{j}$ is a sparse coefficient vector for the $j$-th sample, which means most of the entries in $\mathbf{a}_{j}$ are zero.
The fast selection method based on canonical correlation is performed again, with $\mathbf{D}$ as the target, to find $q$ significant samples from the sample matrix $\mathbf{X}$. For selection methods which evaluate the linear association between candidates, no additional information is gained when the number of selected samples $q$ exceeds the rank of the data matrix $p$. Therefore, to avoid invalid sample selection due to $q>p$, samples are selected in separate batches by the mini-batch sample selection algorithm, which is given in Algorithm 1. Within each batch, the redundancy and interaction of samples are considered, while they are ignored between batches.
Input: $X\in\mathbb{R}^{p\times N}$ (sample matrix), $Y\in\mathbb{R}^{p\times K}$ (target matrix), $q$ (number of samples to select)
Output: $d$ (selected indices)
Step 1: Compute the batch size $b=$ $\lceil q/K\rceil$ and then generate the batch matrix $B\in\mathbb{R}^{K\times s}$, where $B[i,j]\leq b$ and $\sum B=q$;
Step 2: Initialize the candidate sample matrix $X_{c}\leftarrow X$;
for $i\leftarrow 1$ to $K$ do
       Let $y_{i}\leftarrow Y[:,i]$;
       for $j\leftarrow 1$ to $s$ do
             for $k\leftarrow 1$ to $B[i,j]$ do
                   Compute the MCC between each column of $X_{c}$ and $y_{i}$;
                   Select the index $l^{*}$ corresponding to the largest MCC;
                   Append $l^{*}$ to $d$;
                   Remove $X_{c}[:,l^{*}]$ from the candidate matrix;
return $d$;
Algorithm 1 Mini-batch Sample Selection
@@ 3 Numerical Case Studis
3.1 Input Data Analysis
Data pruning is crucial in system identification for the following reasons:
  1. 1.
    Firstly, system identification typically involves time-series data sampled from continuous physical processes. When the sampling interval is small, consecutive data points are highly similar, leading to a strong temporal correlation and redundancy. The larger dataset size due to redundant data increases the computational cost of model training.
  2. 2.
    Secondly, the time-shift operation, commonly employed in system identification introduces delayed versions of the signals as new features, which makes the redundancy of the dataset worse. Redundant features do not provide additional information but increase the complexity of the model, making it more prone to overfitting and harder to interpret.
@@ 3.1 Input Data Analysis
The feature-wise redundancy is illustrated in the correlation heatmap of Fig. 3, where the $i^{th}$ feature corresponds to $y(t-i\Delta t)$. The high Pearson’s correlation near the main diagonal indicates redundancy between neighbouring features, such as $y(t-i\Delta t)$ and $y(t-(i+1)\Delta t)$. This kind of redundancy can be mitigated via feature selection.
Fig. 4 shows the redundancy within the time series samples using principal component analysis (PCA). After projecting the 21-dimensional data into two-dimensional space, it is found that the data forms a continuous trajectory over time. The proximity of adjacent samples in the lower-dimensional space given by PCA indicates that there is redundancy between these time series samples. This redundancy can be addressed by data pruning.
Figure 3: Correlation heatmap illustrating feature-wise redundancy
Figure 4: Correlation heatmap illustrating sample-wise redundancy

### v2

@@ Abstract
System identification is normally involved in augmenting time series data by time shifting and nonlinearisation (e.g., polynomial basis), both of which introduce redundancy in features and samples. Many research works focus on reducing redundancy feature-wise, while less attention is paid to sample-wise redundancy. This paper proposes a novel data pruning method, called mini-batch FastCan, to reduce sample-wise redundancy based on dictionary learning. Time series data is represented by some representative samples, called atoms, via dictionary learning. The useful samples are selected based on their correlation with the atoms. The method is tested on one simulated dataset and two benchmark datasets. The R-squared between the coefficients of models trained on the full datasets and the coefficients of models trained on pruned datasets is adopted to evaluate the performance of data pruning methods. It is found that the proposed method significantly outperforms the random pruning method.
@@
Keywords Data pruning, system identification, dictionary learning, NARX
@@ 1 Introduction
To avoid model overfitting and reduce computational complexity, feature selection methods based on orthogonalisation techniques and greedy search are widely applied in constructing NARX models [10, 11, 12]. The orthogonalisation-based method was firstly derived in [10] to efficiently decide which terms should be included in the Nonlinear Autoregressive Moving Average with eXogenous input models. This method was then further developed in [11] as orthogonal forward-regression estimators to identify parsimonious models of structure-unknown systems by modifying and augmenting some orthogonal least squares methods. A more comprehensive review of this feature selection idea and its development for system identification can be found in [12].
@@
Similar to feature selection, input selection is also a crucial step in system identification, which can help to improve the identification performance and reduce storage and training costs [13, 14, 12]. While optimal selection of input samples (often referred to as data pruning) is currently an active research topic for computer vision [15, 16, 17] and natural language processing [18, 19], its application to system identification has received less attention, although it remains an intractable task. Data pruning plays a vital role in system identification for the following reasons. First, time-series data collected from continuous physical processes often exhibit strong temporal correlation, especially when sampled at high frequencies, resulting in redundant observations that hinder efficient model training. Second, the common use of time-shift operations to generate delayed signal versions further exacerbates data redundancy.
@@
This paper proposes a data-pruning method based on dictionary learning to select useful time series samples for constructing a concise mathematical model to represent a system accurately. Specifically, the reduced polynomial NARX method is applied to the identification of nonlinear dynamic systems as an example. It is worth noting that the proposed method can also be used to select informative samples for other types of system identification models. The canonical-correlation-based fast feature selection method introduced in [20, 21] is adopted to find the most important terms that should be included in the NAXR model to achieve the required accuracy. Dictionary learning is then creatively combined with the Fast selection method based on Canonical correlation (FastCan) to find the most useful time series samples to train these terms.
@@
Dictionary learning, initially popular in the signal processing area, is used to learn an overcomplete dictionary composed of signal atoms, where a sparse linear combination of these atoms can generate the original signal [22]. It is currently widely studied for image processing since images usually admit sparse representation [23, 24]. For the proposed method, k-means-based dictionary learning, introduced in Section 2.2, provides a dictionary of time terms, which is used as pseudo-labels in the data pruning process.
@@
In this paper, the proposed method, named mini-batch FastCan, is compared with the random pruning method. Random pruning commonly serves as a reliable benchmark in sample selection tasks and often yields better results than more advanced methods when 30% or less of the data is retained [25]. The reduced NARX trained with the full dataset is used as the baseline model. The coefficients learned with the pruned dataset is compared with the coefficients of the baseline model. The more the coefficients are close to the baseline model, the better the pruned dataset is. The schematic diagram of the research in this paper is illustrated in Fig. 1.
@@
The next section introduces the details of the proposed data-pruning method for system identification, followed by numerical case studies to demonstrate the necessity and advantages of the method. Subsequently, case studies on two public benchmark datasets [26, 27] for nonlinear system identification are presented, with a discussion on the effect of hyperparamteres in the proposed method.
@@
Figure 1: The schematic diagram of this research. The time series $y(t)$ is generated by the target equation, then time-shifted and nonlinearised to construct a redundant term library. Feature-wise redundancy is represented by a correlation heatmap, while sample-wise redundancy is visualised by a principal component analysis (PCA) plot. Feature selection is applied to the term library to reduce feature-wise redundancy. Dataset A [^1] The colour style used this figure is from bg-mpl-stylesheets [28]., which contains the full set of samples, is used to train the Baseline model. Dataset B, a random subset of Dataset A, is used to train the Random pruned model. Dataset C, consisting of samples selected based on the atoms learned from Dataset A, is used to train the Mini-batch FastCan pruned model. The selected samples obtained by both methods are visualised via a PCA plot. To compare the data pruning methods quantitatively, Dataset A is pruned repeatedly by random selection and mini-batch FastCan, and the models are then trained on the pruned datasets, respectively. Data pruning performance is then evaluated by the R-squared score between the coefficients learned from the pruned samples and the coefficients of the baseline model.
@@ 2 Methodology
This section introduces the dictionary learning-based data pruning method, mini-batch FastCan, for a discrete system with finite orders.
@@ 2.1 Reduced Polynomial NARX
The reduced polynomial NARX model is a nonlinear dynamic model that represents the system output as a sparse subset of polynomial terms consisting of past outputs and inputs selected from the full NARX model structure. It reduces complexity by including only important or relevant terms, rather than all possible combinations. Mathematically, a NARX model is used to simulate a system, which is formulated as
$\begin{split}y(k)=&F[y(k-1),y(k-2),...,y(k-n_{y}),\\ &u(k-1),u(k-2),...,u(k-n_{u})]+e(k)\end{split}$ (1)
where $y(k)$, $u(k)$ and $e(k)$ are the system output, input and noise sequences, respectively. $n_{y}$ and $n_{u}$ are the maximum lags for the system output and input. $F[\cdot]$ is a nonlinear function.
@@
$\begin{split}y(k)=\theta_{0}&+\sum_{i_{1}=1}^{n}f_{i_{1}}\big{(}x_{i_{1}}(k)\big{)}\\ &+\sum_{i_{1}=1}^{n}\sum_{i_{2}=i_{1}}^{n}f_{i_{1}i_{2}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k)\big{)}+\dots\\ &+\sum_{i_{1}=1}^{n}\cdots\sum_{i_{\ell}=i_{\ell-1}}^{n}f_{i_{1}i_{2}\cdots i_{\ell}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k),\dots,x_{i_{\ell}}(k)\big{)}+e(k)\end{split}$ (2)
@@
$f_{i_{1}i_{2}\cdots i_{d}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k),\dots,x_{i_{d}}(k)\big{)}=\theta_{i_{1}i_{2}\cdots i_{d}}\prod_{j=1}^{d}x_{i_{j}}(k),\quad$ $1\leq d\leq\ell$
$x_{i_{j}}(k)=\begin{cases}y(k-i),&1\leq i\leq n_{y}\\ u(k-i+n_{y}),&n_{y}+1\leq i\leq n=n_{y}+n_{u}\end{cases}$
where $\theta_{i_{1}i_{2}\cdots i_{d}}$ are model parameters, and $\prod_{j=1}^{d}x_{i_{j}}(k)$ are model terms whose order is not higher than $\ell$.
The total number of model terms in the polynomial NARX model given in equation (2) is $M=(n+\ell)!/[n!\,\ell!]$. It can be seen that the full NARX model can include a large number of terms, increasing the risk of overfitting. Because only a subset of these terms is typically important to capture the underlying dynamic relationship [2], the canonical-correlation-based fast feature selection is carried out here to find out the $m$ significant model terms from $\prod_{j=1}^{d}x_{i_{j}}(k)$ with $y(k)$ as the target. Therefore, the reduced polynomial NARX model can be given as,
$y(k)=\theta_{0}+\sum_{j=1}^{m}\sum_{i_{1}=1}^{n}\cdots\sum_{i_{d}=i_{d-1}}^{n}f_{i_{1}i_{2}\cdots i_{d}}\big{(}x_{i_{1}}(k),x_{i_{2}}(k),\dots,x_{i_{d}}(k)\big{)}+e(k),\quad 1\leq m\leq M$
Refer to [20, 21] for details on the specific feature selection steps.
2.2 Data Pruning with Mini-batch FastCan
@@
$\mathbf{X}=\begin{pmatrix}x_{1,1}&\dots&x_{1,N}\\ \vdots&\ddots&\vdots\\ x_{m,1}&\dots&x_{m,N}\end{pmatrix}$
where $m$ is the number of the features, i.e., the selected model terms, and $N$ is the number of time series samples given these selected features.
A dictionary $\mathbf{D}\in\mathbb{R}^{m\times q}$ for the matrix $\mathbf{X}\in\mathbb{R}^{m\times N}$ is obtained by mini-batch k-means clustering [29] and used as the target matrix for the subsequent data pruning step, where $q$ is the number of atoms in the dictionary, and typically $q>m$ for an overcomplete dictionary. Each data sample $\mathbf{x}_{j}\in\mathbb{R}^{m}$ (column of $\mathbf{X}$) can be approximated as a sparse linear combination of the dictionary atoms, given by,
@@
where $\mathbf{a}_{j}\in\mathbb{R}^{q}$ is a sparse coefficient vector for the $j$-th sample, which means most of the entries in $\mathbf{a}_{j}$ are zero.
Input: $\mathbf{X}\in\mathbb{R}^{m\times N}$ ;
  $\triangleright$ Sample matrix
$q\in\mathbb{N}$ ;
  $\triangleright$ Number of atoms in a dictionary $\mathbf{D}$
$p\in\mathbb{N}$ ;
  $\triangleright$ Batch size; Optional
$n\in\mathbb{N}$ ;
  $\triangleright$ Number of samples to select
Output: $\mathbf{s}\in\mathbb{R}^{1\times n}$ ;
  $\triangleright$ Selected indices
Step 1:
Apply the k-means-based dictionary learning [29] to $\mathbf{X}^{\top}$ with $q$ clusters;
The resulting $q$ cluster centres form the columns of a dictionary $\mathbf{D}\in\mathbb{R}^{m\times q}$ ;
  $\triangleright$ Target matrix
Step 2:
if $p$ is not specified or $p>\lceil n/q\rceil$ then
       Set $p\leftarrow\lceil n/q\rceil$;
      if $p>m$ then
             Set $p\leftarrow m$;
Step 3:
Generate the batch matrix $\mathbf{B}\in\mathbb{R}^{q\times t}$, where $t=$ $\lceil n/(q\times p)\rceil$, $\mathbf{B}[i,j]$ $\leq p$ and $\sum_{i=1}^{q}\sum_{j=1}^{t}\mathbf{B}[i,j]=n$;
Step 4:
Initialize the candidate sample matrix $\mathbf{X}_{c}\leftarrow\mathbf{X}$ and the target matrix $\mathbf{D}_{c}\leftarrow\mathbf{D}$;
for $i\leftarrow 1$ to $q$ do
       Let $\mathbf{d}_{i}\leftarrow\mathbf{D}_{c}[:,i]$;
       for $j\leftarrow 1$ to $t$ do
             Select $\mathbf{B}[i,j]$ samples from $\mathbf{X}_{c}$ by using the canonical-correlation-based fast selection method [21], with $\mathbf{d}_{i}$ serving as the target vector;
             Append the indices of the selected samples to $\mathbf{s}$;
             Remove the selected samples from the candidate matrix $\mathbf{X}_{c}$;
return $\mathbf{s}$;
Algorithm 1 Mini-batch FastCan
The canonical-correlation-based fast selection method [21] is performed again, with $\mathbf{D}$ as the target, to find $n$ significant samples from the sample matrix $\mathbf{X}$. For selection methods which evaluate the linear association between candidates, no additional information is gained once the number of selected samples $n$ exceeds the rank of the data matrix $m$. Therefore, to avoid invalid sample selection due to $n>m$, samples are selected in separate batches by the mini-batch FastCan method, whose pseudocode is given in Algorithm 1, and the corresponding codes are available in the GitHub repository  [^2] https://github.com/MatthewSZhang/data-pruning-sysid [^3] https://github.com/scikit-learn-contrib/fastcan. Within each batch, the redundancy and interaction between samples are considered, while these are ignored between batches.
In Algorithm 1, if the batch size $p$ is not specified, it is set to $\lceil q/k\rceil$. If a given $p>\lceil n/q\rceil$, it is also reset to $\lceil q/k\rceil$, since exceeding this threshold will result in some atoms being excluded from the selection process. Additionally, this algorithm allows the number of atoms $q$ to be smaller than the number of features $m$, enabling a larger batch size to better capture sample redundancy.
@@ 3 Numerical Case Studis
3.1 Visualisation of Sample Redundancy in System Identification
@@
$t$ (s) $y(t-\Delta t)$ $y(t-2\Delta t)$ $y(t-3\Delta t)$ $y(t-20\Delta t)$
0 NaN NaN NaN NaN
0.01 0.000 NaN NaN NaN
0.02 0.063 0.000 NaN NaN
0.03 0.127 0.063 0.000 NaN
0.04 0.189 0.127 0.063 NaN
0.99 -0.063 -0.127 -0.189 -0.955
(a) feature-wise
(b) sample-wise
Figure 3: Correlation heatmap illustrating redundancy in data
The feature-wise redundancy is illustrated in the correlation heatmap of Fig. 3a, where the $i^{th}$ feature corresponds to $y(t-i\Delta t)$ with $1\leq i\leq 20$. The high Pearson’s correlation near the main diagonal indicates redundancy between neighbouring features, such as $y(t-i\Delta t)$ and $y(t-(i+1)\Delta t)$. This kind of redundancy can be mitigated via feature selection.
Fig. 3b shows the redundancy within the time series samples using principal component analysis (PCA). After projecting the 20-dimensional data into two-dimensional space, it is found that the data forms a continuous trajectory over time. The proximity of adjacent samples in the lower-dimensional space given by PCA indicates that there is redundancy between these time series samples. This redundancy can be addressed by data pruning.